# Neural Topic Modeling with large language models in the loop

### v1

@@ 6 Conclusion
In this paper, we introduced LLM-ITL, a novel framework that integrates Large Language Models (LLMs) with Neural Topic Models (NTMs) to address the limitations of both traditional topic models and the direct use of LLMs for topic discovery. By incorporating a confidence-weighted Optimal Transport (OT)-based topic alignment, LLM-ITL improves the interpretability and coherence of topics while maintaining the quality of document representations. Our framework effectively leverages the strengths of both LLMs and NTMs, offering a flexible, scalable, and efficient solution for topic modeling. Extensive experiments on benchmark datasets demonstrate that the LLM-ITL variants of NTMs achieve state-of-the-art performance in terms of topic coherence and document representation. In terms of limitations, the framework’s reliance on LLM-generated refinements introduces a dependency on the quality of the LLM’s outputs, which may vary based on the model used. Moreover, the method has been primarily evaluated on benchmark datasets, and its performance in more domain-specific corpora may require further investigation.
@@ Appendix A Prompts
A.1 Topic suggestion
Figure A1: Prompt and output of topic suggestion with CoT. We take the product of token probabilities of topic label (e.g., words in red color) as the label token probability. We take the proportion of intruders (e.g., words in green color) as the word intrusion confidence.
A.2 Topic suggestion with verbalized confidence
Figure A2: Prompt and output of topic suggestion with CoT and verbalized confidence. The number in red color in the LLM’s output represents the verbalized confidence.
A.3 Self-reflect confidence
@@ Appendix B Algorithm
Algorithm 1 Algorithm for LLM-ITL
Input:
Train documents; An LLM; Pre-trained word embeddings; Hyper-parameters $T^{\text{refine}}$, $\gamma$; Training iteration $I$; Number of topics $K$.
@@ Appendix C Related work: LLM Uncertainty Estimation
Uncertainty estimation for LLMs (Geng et al., 2024) is emerging with the rapid usage of LLMs and their risk of hallucination (Ji et al., 2023). Sequence probability (Ren et al., 2022) leverages token probabilities to measure answer confidence. Verbalized confidence (Tian et al., 2023; Xiong et al., 2023) utilizes the LLM’s own capability to evaluate its answer uncertainty. Consistency-based (Lin et al., 2023; Manakul et al., 2023) approaches sample multiple outputs from the LLM and measure answer consistency as uncertainty. Entropy-based (Kuhn et al., 2023; Hou et al., 2023) approaches use multiple LLM outputs to estimate the output space and compute entropy as uncertainty for the answer. Hybrid frameworks (Chen & Mueller, 2023; Gao et al., 2024) combine different approaches for a comprehensive estimation. Internal states (Chen et al., 2024) are another useful source for LLM uncertainty quantification. Unlike those works that estimate uncertainty for LLMs in general natural language generation tasks, ours focuses on task-specific uncertainty of the LLM in suggesting topic words.
@@ D.1 Details of Dataset
We conduct experiments on 20News  [^12] https://huggingface.co/datasets/SetFit/20_newsgroups, R8  [^13] https://huggingface.co/datasets/yangwang825/reuters-21578, DBpedia  [^14] https://huggingface.co/datasets/fancyzhx/dbpedia_14, and AGNews  [^15] https://huggingface.co/datasets/fancyzhx/ag_news. For DBpedia and AGNews, we randomly sample a subset of 20,000 documents. We retain the original text documents for models that accept text as input, and preprocess the documents into Bag-of-Words (BOW) format for models that are trained on BOWs. We convert the documents into BOW vectors through the following steps: First, we clean the documents by removing special characters and stop words, followed by tokenization. Next, we build the vocabulary by including words with a document frequency greater than five and less than 80% of the total documents. Since we use the pre-trained word embeddings of GloVe (Pennington et al., 2014), we further filter the vocabulary by retaining only the words that are in the GloVe vocabulary. Finally, we transform the documents into BOWs based on the filtered vocabulary set. The statistics of the preprocessed datasets are summarized in Table D1.
@@ D.2 Details of Baselines
We run the following topic models as our baselines, including Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that generates documents by mixtures of topics; Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneering NTM based on the VAE framework; LDA with Products of Experts (PLDA) (Srivastava & Sutton, 2017), an NTM that uses a product of experts instead of the mixture model in LDA; Embedded Topic Model (ETM) (Dieng et al., 2020), which involves word and topic embeddings in the generative process of documents; Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2017), an NTM that leverages extra information from metadata; Contrastive Learning Neural Topic Model (CLNTM) (Nguyen & Luu, 2021), an NTM that is based on the contrastive learning framework; BERTopic (Grootendorst, 2022), a recent clustering-based topic model that utilizes embeddings from pre-trained language models; TopicGPT (Pham et al., 2023), a latest LLM-based topic model that leverages an LLM for topic generation and assignment.
As for the implementations of baseline models, we use Mallet  [^16] https://radimrehurek.com/gensim_3.8.3/models/wrappers/ldamallet.html for LDA with Gibbs sampling, and the original implementations for the other models. For NTMs including NVDM, PLDA, ETM, SCHOLAR and CLNTM, we tune their hyper-parameters for our datasets; For BERTopic, we fine-tune the topic representations after the topics are learned, as suggested by their implementation [^17] https://maartengr.github.io/BERTopic/index.html. For TopicGPT, we use GPT-4 for topic generation and GPT-3.5 for topic assignment, randomly sampling 600 documents from the training set for each dataset, as suggested by their paper. We run all models except TopicGPT five times in each experiment and report the mean and standard deviation of their performance. For TopicGPT, we run it once for each experiment using a temperature value of zero to enable deterministic output, following the setting of their paper.
@@ D.3 Details of Evaluation Metrics
For topic evaluation, we apply the commonly-used topic coherence metric, Normalized Pointwise Mutual Information (NPMI) (Lau et al., 2014), which evaluates topic coherence based on the co-occurrence of the topic’s top words in a reference corpus. We use Wikipedia as the reference corpus for NPMI and consider the top 10 words of each topic, with implementation done using the Palmetto package [^18] https://github.com/dice-group/Palmetto (Röder et al., 2015). We report the average NPMI score (in percentage) of all learned topics. For documents’ topical representation (i.e., topic proportion) evaluation, a common practice is to compare the document clusters formed by topic proportions with those formed by the documents’ true labels, known as topic alignment. Following previous works (Chuang et al., 2013; Pham et al., 2023), we assign each test document to a cluster based on the top-weighted topic of its topical representation, and compute Purity and Normalized Mutual Information (NMI) based on the documents’ cluster assignments and their true labels. As Purity and NMI are often reported together and within the same range, we report the average score of both metrics as PN. For all evaluations, we use the model state at the end of the training iteration to compute all evaluation metrics.
@@ E.1 Topic coherence at different K
Figure E1: Topic coherence (NPMI) at different settings of the number of topics (i.e., $K$) on (a) 20News, (b) R8, (c) DBpedia and (d) AGNews.
Here, we illustrate the topic coherence performance of topic models across different numbers of topics (i.e., $K$). The results are shown in Figure E1. We observe that LLM-ITL consistently enhances topic coherence in its base models, and achieves state-of-the-art performance across various settings of $K$ in most cases.
@@ E.2 Topic alignment at different K
Figure E2: Topic alignment (PN) at different settings of the number of topics (i.e., $K$) on (a) 20News, (b) R8, (c) DBpedia and (d) AGNews.
Here, we illustrate the topic alignment performance of topic models across different numbers of topics (i.e., $K$). The results are shown in Figure E2. We observe that LLM-ITL consistently inherits the topic alignment performance of its base model across different settings of $K$. The performance drop with an increase in $K$ for SCHOLAR and CLNTM in short document datasets (e.g., DBpedia and AGNews) is due to their sensitivity to $K$ in short documents.
E.3 Topic diversity at different K
Figure E3: Topic diversity (TD) at different settings of the number of topics (i.e., $K$) on (a) 20News, (b) R8, (c) DBpedia and (d) AGNews.
Here, we illustrate the topic diversity (TD) performance of topic models across different numbers of topics (i.e., $K$). The results are presented in Figure E3. We observe that LLM-ITL demonstrates comparable performance in terms of topic diversity compared to its base models in most cases.
E.4 Purity & NMI
As we report PN (the mean of Purity and NMI) as an overall topic alignment metric in previous sections, we illustrate detailed Purity and NMI performance in this section. As shown in Table E1 and E2, Figure E4 and E5, LLM-ITL consistently inherits topic alignment performance from its base models in terms of both Purity and NMI, which is consistent with our previous observations based on PN.
Table E1: Purity performance. The best and second-best scores are highlighted in boldface and underlined, respectively. The performance improvement of LLM-ITL over its base model is computed.
@@ E.4 Purity & NMI
LDA 0.521 ± 0.012 0.920 ± 0.004 0.813 ± 0.018 0.818 ± 0.010
NVDM 0.169 ± 0.007 0.603 ± 0.013 0.220 ± 0.004 0.724 ± 0.013
PLDA 0.130 ± 0.004 0.771 ± 0.011 0.730 ± 0.007 0.724 ± 0.013
BERTopic 0.371 ± 0.010 0.875 ± 0.002 0.748 ± 0.009 0.648 ± 0.014
TopicGPT 0.336 ± 0.000 0.577 ± 0.000 0.718 ± 0.000 0.819 ± 0.000
ETM 0.410 ± 0.011 0.875 ± 0.005 0.794 ± 0.016 0.784 ± 0.008
LLM-ITL (ETM) 0.403 ± 0.015 0.875 ± 0.012 0.761 ± 0.019 0.784 ± 0.006
$\downarrow$ 0.007 $\uparrow$ $0$downarrow 0.033 $\uparrow$ 0.000
SCHOLAR 0.627 ± 0.014 0.911 ± 0.015 0.872 ± 0.018 0.855 ± 0.004
LLM-ITL (SCHOLAR) 0.607 ± 0.014 0.911 ± 0.015 0.875 ± 0.015 0.855 ± 0.004
$\downarrow$ 0.020 $\uparrow$ $0$uparrow 0.003 $\uparrow$ 0.000
CLNTM 0.623 ± 0.015 0.923 ± 0.009 0.725 ± 0.047 0.821 ± 0.013
LLM-ITL (CLNTM) 0.623 ± 0.006 0.923 ± 0.008 0.727 ± 0.045 0.806 ± 0.014
$\uparrow$ $0$uparrow $0$uparrow 0.002 $\downarrow$ 0.015
Table E2: NMI performance. The best and second-best scores are highlighted in boldface and underlined, respectively. The performance improvement of LLM-ITL over its base model is computed.

### v2

@@ 6 Conclusion
In this paper, we introduced LLM-ITL, a novel framework that integrates Large Language Models (LLMs) with Neural Topic Models (NTMs) to address the limitations of both traditional topic models and the direct use of LLMs for topic discovery. By incorporating a confidence-weighted Optimal Transport (OT)-based topic alignment, LLM-ITL improves the interpretability and coherence of topics while maintaining the quality of document representations. Our framework effectively leverages the strengths of both LLMs and NTMs, offering a flexible, scalable, and efficient solution for topic modeling. Extensive experiments on benchmark datasets demonstrate that applying LLM-ITL with NTMs significantly boosts topic interpretability while maintaining document representation quality.
7 Limitation
The proposed framework relies on the refinements generated by the LLM. Over-reliance on the LLM during the refinement process may introduce out-of-corpus information or bias the topics toward the LLM’s pre-training knowledge. This can negatively impact the alignment of learned topics with the input corpus, particularly in cases where the corpus diverges significantly from the LLM’s training data.
@@ Appendix A Prompts
Figure A1: Prompt and output of topic suggestion with CoT. We take the product of token probabilities of topic label (e.g., words in red color) as the label token probability. We take the proportion of intruders (e.g., words in green color) as the word intrusion confidence. $N=2$ and $M=10$ in this example.
Figure A2: Prompt and output of topic suggestion with CoT and verbalized confidence. The number in red color in the LLM’s output represents the verbalized confidence. $N=2$ and $M=10$ in this example.
@@ Appendix B Algorithm
Input: Train documents; An LLM; Pre-trained word embeddings; Hyper-parameters $T^{\text{refine}}$, $\gamma$; Training iteration $I$; Number of topics $K$.
@@ Appendix C Related work: LLM Uncertainty Estimation
Uncertainty estimation for LLMs (Geng et al., 2024) is emerging with the rapid usage of LLMs and their risk of hallucination (Ji et al., 2023). Sequence probability (Ren et al., 2022) leverages token probabilities to measure answer confidence. Verbalized confidence (Tian et al., 2023; Xiong et al., 2023) utilizes the LLM’s own capability to evaluate its answer uncertainty. Consistency-based (Lin et al., 2023; Manakul et al., 2023) approaches sample multiple outputs from the LLM and measure answer consistency as uncertainty. Entropy-based (Kuhn et al., 2023; Hou et al., 2023) approaches estimate the output space from multiple LLM outputs and compute entropy as uncertainty for the answer. Hybrid frameworks (Chen and Mueller, 2023; Gao et al., 2024) combine different approaches for a comprehensive estimation. Internal states (Chen et al., 2024) are another useful source for LLM uncertainty quantification. Unlike those works that estimate uncertainty for LLMs in general natural language generation tasks, ours focuses on task-specific uncertainty of the LLM in suggesting topic words.
@@ D.1 Details of Dataset
Table D1: Statistics of the Datasets
We conduct experiments on 20News [^3] https://huggingface.co/datasets/SetFit/20_newsgroups, R8444https://huggingface.co/datasets/yangwang825/reuters-21578, DBpedia [^5] https://huggingface.co/datasets/fancyzhx/dbpedia_14, and AGNews [^6] https://huggingface.co/datasets/fancyzhx/ag_news. For DBpedia and AGNews, we randomly sample a subset of 20,000 documents. We retain the original text documents for models that accept text as input, and preprocess the documents into Bag-of-Words (BOW) format for models that are trained on BOWs. We convert the documents into BOW vectors through the following steps: First, we clean the documents by removing special characters and stop words, followed by tokenization. Next, we build the vocabulary by including words with a document frequency greater than five and less than 80% of the total documents. Since we use the pre-trained word embeddings of GloVe (Pennington et al., 2014), we further filter the vocabulary by retaining only the words that are in the GloVe vocabulary. Finally, we transform the documents into BOWs based on the filtered vocabulary set. The statistics of the preprocessed datasets are summarized in Table D1.
@@ D.2 Details of Baselines
We run the following topic models in our experiments, including Latent Dirichlet Allocation (LDA) (Blei et al., 2003), the most popular probabilistic topic model that generates documents by mixtures of topics; Neural Variational Document Model (NVDM) (Miao et al., 2017), a pioneering NTM based on the VAE framework; LDA with Products of Experts (PLDA) (Srivastava and Sutton, 2017), an NTM that uses a product of experts instead of the mixture model in LDA; Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2017), an NTM that leverages extra information from metadata; Embedded Topic Model (ETM) (Dieng et al., 2020), which involves word and topic embeddings in the generative process of documents; Neural Sinkhorn Topic Model (NSTM) (Zhao et al., 2020), an NTM based on an optimal transport framework; Contrastive Learning Neural Topic Model (CLNTM) (Nguyen and Luu, 2021), an NTM that is based on the contrastive learning framework; BERTopic (Grootendorst, 2022), a recent clustering-based topic model that utilizes embeddings from pre-trained language models; WeTe (Wang et al., 2022), which represents a mixture of word embeddings as a mixture of topic embeddings; Embedding Clustering Regularization Topic Model (ECRTM) (Wu et al., 2023), an NTM with embedding clustering regularization; TopicGPT (Pham et al., 2023), a latest LLM-based topic model that leverages an LLM for topic generation and assignment.
As for the implementations of baseline models, we use Mallet [^7] https://radimrehurek.com/gensim_3.8.3/models/wrappers/ ldamallet.html for LDA with Gibbs sampling, and the original implementations for the other models. For NTMs including NVDM, PLDA, SCHOLAR, ETM, NSTM, CLNTM, WeTe and ECRTM, we tune their hyper-parameters for the datasets; For BERTopic, we fine-tune the topic representations after the topics are learned, as suggested by their implementation [^8] https://maartengr.github.io/BERTopic/index.html. For TopicGPT, we use GPT-4 for topic generation and GPT-3.5 for topic assignment, randomly sampling 600 documents from the training set for each dataset, as suggested by their paper. We run all models except TopicGPT five times in each experiment and report the mean and standard deviation of their performance. For TopicGPT, we run it once for each experiment using a temperature value of zero to enable deterministic output, following the setting of their paper.
@@ D.3 Details of Evaluation Metrics
For topic evaluation, we apply the commonly-used topic coherence metric, $\bm{C_{V}}$ (Röder et al., 2015), which evaluates topic coherence based on the co-occurrence of the topic’s top words in a reference corpus. Following standard evaluation protocol, we use Wikipedia as the reference corpus and consider the top 10 words of each topic, with implementation done using the Palmetto package [^9] https://github.com/dice-group/Palmetto (Röder et al., 2015). We report the average $C_{V}$ score of all learned topics as the overall topic coherence performance. For documents’ topical representation (i.e., topic proportion) evaluation, a common practice is to compare the document clusters formed by topic proportions with those formed by the documents’ true labels, known as topic alignment. Following previous works (Chuang et al., 2013; Pham et al., 2023), we assign each test document to a cluster based on the top-weighted topic of its topical representation, and compute Purity and Normalized Mutual Information (NMI) based on the documents’ cluster assignments and their true labels. As Purity and NMI are often reported together and within the same range, we report the average score of both metrics as PN. For all evaluations, we use the model state at the end of the training iteration to compute all evaluation metrics.
@@ E.1 Topic Coherence at Different K
We illustrate the topic coherence performance ($C_{V}$) of topic models across different numbers of topics (i.e., $K$) in Table E1. Overall, using LLM-ITL significantly boosts topic coherence across all cases in our experiments, with minimum gains of +10.4% (NSTM as the base model on DBpedia at $K=50$) and maximum gains of +86.3% (ECRTM as the base model on 20News at $K=25$) over the base model.
@@ E.2 Topic Alignment at Different K
We illustrate the topic alignment performance (PN) of topic models across different numbers of topics (i.e., $K$) in Table E2. Overall, using LLM-ITL inherits the topic alignment performance of its base model, with slight changes ranging from -5.2% (NSTM as the base model on 20News at $K=100$) to +5.2% (PLDA as the base model on 20News at $K=75$).
E.3 Topic Diversity (TD)
We evaluate the performance of topic diversity (TD) defined by Dieng et al. (2020), which evaluates the distinctiveness of the top-words within the learned topics. The results across different numbers of topics (i.e., $K$) is illustrated in Table E3. Overall, applying LLM-ITL results in changes to TD ranging from -34.6% (ECRTM as the base model on 20News at $K=100$) to +55.5% (NSTM as the base model on 20News at $K=100$) compared to its base model. A further analysis of overall topic quality, which considers both topic coherence and diversity, is described in the following section.
E.4 Overall Topic Quality (TQ)
[…]
E.5 Purity & NMI
As we report PN (the mean of Purity and NMI) as an overall topic alignment metric in previous sections, here we present the detailed performance of Purity and NMI in Table E5 and Table E6, respectively.
@@ E.5 Purity & NMI
K=25 K=75 K=100 K=25 K=75 K=100 K=50 K=75 K=100 K=50 K=75 K=100
LDA 0.558 ± 0.008 0.525 ± 0.005 0.510 ± 0.004 0.387 ± 0.011 0.373 ± 0.003 0.374 ± 0.006 0.529 ± 0.005 0.520 ± 0.010 0.506 ± 0.004 0.526 ± 0.010 0.502 ± 0.003 0.480 ± 0.003
BERTopic 0.435 ± 0.029 0.403 ± 0.012 0.390 ± 0.005 0.349 ± 0.013 0.329 ± 0.008 0.333 ± 0.007 0.559 ± 0.011 0.537 ± 0.007 0.516 ± 0.006 0.500 ± 0.016 0.485 ± 0.004 0.479 ± 0.004
TopicGPT NA NA NA NA NA NA NA NA NA NA NA NA
NVDM 0.299 ± 0.012 0.241 ± 0.004 0.232 ± 0.004 0.324 ± 0.011 0.283 ± 0.002 0.270 ± 0.002 0.274 ± 0.003 0.260 ± 0.004 0.259 ± 0.005 0.289 ± 0.004 0.266 ± 0.005 0.256 ± 0.007
+ LLM-ITL 0.439 ± 0.019 0.298 ± 0.009 0.292 ± 0.004 0.438 ± 0.017 0.358 ± 0.011 0.340 ± 0.006 0.365 ± 0.008 0.353 ± 0.008 0.346 ± 0.006 0.357 ± 0.018 0.326 ± 0.005 0.325 ± 0.005
$\uparrow$ 46.8% $\uparrow$ 23.7% $\uparrow$ 25.9% $\uparrow$ 35.2% $\uparrow$ 26.5% $\uparrow$ 25.9% $\uparrow$ 33.2% $\uparrow$ 35.8% $\uparrow$ 33.6% $\uparrow$ 23.5% $\uparrow$ 22.6% $\uparrow$ 27.0%
PLDA 0.418 ± 0.020 0.352 ± 0.010 0.336 ± 0.002 0.391 ± 0.008 0.365 ± 0.006 0.348 ± 0.007 0.484 ± 0.008 0.457 ± 0.009 0.438 ± 0.004 0.515 ± 0.007 0.490 ± 0.012 0.470 ± 0.003
+ LLM-ITL 0.557 ± 0.026 0.534 ± 0.010 0.506 ± 0.013 0.477 ± 0.036 0.456 ± 0.011 0.452 ± 0.004 0.598 ± 0.009 0.570 ± 0.021 0.568 ± 0.015 0.595 ± 0.012 0.582 ± 0.012 0.581 ± 0.008
$\uparrow$ 33.3% $\uparrow$ 51.7% $\uparrow$ 50.6% $\uparrow$ 22.0% $\uparrow$ 24.9% $\uparrow$ 29.9% $\uparrow$ 23.6% $\uparrow$ 24.7% $\uparrow$ 29.7% $\uparrow$ 15.5% $\uparrow$ 18.8% $\uparrow$ 23.6%
SCHOLAR 0.520 ± 0.018 0.478 ± 0.009 0.472 ± 0.014 0.399 ± 0.010 0.385 ± 0.005 0.376 ± 0.004 0.434 ± 0.007 0.422 ± 0.015 0.403 ± 0.021 0.577 ± 0.025 0.581 ± 0.010 0.541 ± 0.014
+ LLM-ITL 0.622 ± 0.028 0.584 ± 0.008 0.562 ± 0.011 0.462 ± 0.010 0.437 ± 0.006 0.424 ± 0.009 0.564 ± 0.037 0.565 ± 0.024 0.556 ± 0.020 0.643 ± 0.015 0.647 ± 0.005 0.627 ± 0.014
$\uparrow$ 19.6% $\uparrow$ 22.2% $\uparrow$ 19.1% $\uparrow$ 15.8% $\uparrow$ 13.5% $\uparrow$ 12.8% $\uparrow$ 30.0% $\uparrow$ 33.9% $\uparrow$ 38.0% $\uparrow$ 11.4% $\uparrow$ 11.4% $\uparrow$ 15.9%
ETM 0.511 ± 0.007 0.466 ± 0.005 0.461 ± 0.008 0.447 ± 0.017 0.432 ± 0.011 0.431 ± 0.009 0.507 ± 0.007 0.501 ± 0.006 0.496 ± 0.007 0.497 ± 0.010 0.497 ± 0.004 0.480 ± 0.004
+ LLM-ITL 0.627 ± 0.019 0.566 ± 0.007 0.553 ± 0.014 0.570 ± 0.014 0.565 ± 0.010 0.580 ± 0.004 0.657 ± 0.031 0.631 ± 0.014 0.626 ± 0.009 0.626 ± 0.018 0.620 ± 0.012 0.618 ± 0.006
$\uparrow$ 22.7% $\uparrow$ 21.5% $\uparrow$ 20.0% $\uparrow$ 27.5% $\uparrow$ 30.8% $\uparrow$ 34.6% $\uparrow$ 29.6% $\uparrow$ 25.9% $\uparrow$ 26.2% $\uparrow$ 26.0% $\uparrow$ 24.7% $\uparrow$ 28.8%
NSTM 0.450 ± 0.008 0.425 ± 0.003 0.424 ± 0.008 0.415 ± 0.014 0.407 ± 0.006 0.402 ± 0.013 0.636 ± 0.008 0.621 ± 0.014 0.616 ± 0.006 0.596 ± 0.015 0.583 ± 0.015 0.575 ± 0.018
+ LLM-ITL 0.510 ± 0.027 0.524 ± 0.008 0.518 ± 0.011 0.549 ± 0.011 0.544 ± 0.021 0.534 ± 0.016 0.702 ± 0.009 0.688 ± 0.009 0.690 ± 0.014 0.661 ± 0.021 0.670 ± 0.008 0.649 ± 0.012
$\uparrow$ 13.3% $\uparrow$ 23.3% $\uparrow$ 22.2% $\uparrow$ 32.3% $\uparrow$ 33.7% $\uparrow$ 32.8% $\uparrow$ 10.4% $\uparrow$ 10.8% $\uparrow$ 12.0% $\uparrow$ 10.9% $\uparrow$ 14.9% $\uparrow$ 12.9%
CLNTM 0.501 ± 0.010 0.478 ± 0.007 0.469 ± 0.010 0.387 ± 0.005 0.346 ± 0.004 0.354 ± 0.002 0.414 ± 0.030 0.394 ± 0.014 0.393 ± 0.011 0.549 ± 0.010 0.567 ± 0.012 0.569 ± 0.010
+ LLM-ITL 0.623 ± 0.011 0.587 ± 0.011 0.580 ± 0.006 0.446 ± 0.006 0.422 ± 0.008 0.411 ± 0.004 0.556 ± 0.004 0.540 ± 0.020 0.547 ± 0.008 0.613 ± 0.003 0.628 ± 0.005 0.658 ± 0.013
$\uparrow$ 24.4% $\uparrow$ 22.8% $\uparrow$ 23.7% $\uparrow$ 15.2% $\uparrow$ 22.0% $\uparrow$ 16.1% $\uparrow$ 34.3% $\uparrow$ 37.1% $\uparrow$ 39.2% $\uparrow$ 11.7% $\uparrow$ 10.8% $\uparrow$ 15.6%
WeTe 0.488 ± 0.010 0.505 ± 0.017 0.498 ± 0.013 0.534 ± 0.008 0.534 ± 0.011 0.533 ± 0.011 0.557 ± 0.012 0.526 ± 0.005 0.516 ± 0.015 0.587 ± 0.007 0.562 ± 0.003 0.556 ± 0.005
+ LLM-ITL 0.539 ± 0.030 0.578 ± 0.022 0.556 ± 0.021 0.604 ± 0.015 0.605 ± 0.012 0.601 ± 0.008 0.650 ± 0.009 0.660 ± 0.011 0.665 ± 0.018 0.655 ± 0.007 0.689 ± 0.010 0.693 ± 0.015
$\uparrow$ 10.5% $\uparrow$ 14.5% $\uparrow$ 11.6% $\uparrow$ 13.1% $\uparrow$ 13.3% $\uparrow$ 12.8% $\uparrow$ 16.7% $\uparrow$ 25.5% $\uparrow$ 28.9% $\uparrow$ 11.6% $\uparrow$ 22.6% $\uparrow$ 24.6%
ECRTM 0.350 ± 0.013 0.324 ± 0.007 0.313 ± 0.003 0.279 ± 0.015 0.287 ± 0.013 0.288 ± 0.009 0.520 ± 0.016 0.436 ± 0.014 0.407 ± 0.008 0.408 ± 0.026 0.382 ± 0.011 0.352 ± 0.014
+ LLM-ITL 0.652 ± 0.026 0.518 ± 0.017 0.520 ± 0.023 0.352 ± 0.012 0.354 ± 0.008 0.362 ± 0.008 0.630 ± 0.006 0.563 ± 0.010 0.521 ± 0.010 0.475 ± 0.019 0.451 ± 0.021 0.426 ± 0.012
$\uparrow$ 86.3% $\uparrow$ 59.9% $\uparrow$ 66.1% $\uparrow$ 26.2% $\uparrow$ 23.3% $\uparrow$ 25.7% $\uparrow$ 21.2% $\uparrow$ 29.1% $\uparrow$ 28.0% $\uparrow$ 16.4% $\uparrow$ 18.1% $\uparrow$ 21.0%
Table E1: Topic coherence ($C_{V}$) at other settings of $K$.